OpenStudy (anonymous):

For the function g(x)= -2(x-1)^2 +7 How do I define the equation of g^-1 by switching the operations

OpenStudy (anonymous):

g(x) = -2(x-1)^2 + 7 to find the inverse g^-1 we swap the x and the y such that y=g^-1(x) x = -2(y -1)^2 + 7 solving for y: x = -2(y -1)^2 + 7 -(x - 7)/2 = (y - 1)^2 y - 1 = +- sqrt((7 - x)/2) y = 1 +- sqrt((7 - x)/2) thus g^-1(x) = 1 +- sqrt((7 - x)/2)

OpenStudy (anonymous):

wait so you brought the -2 to the other side?

OpenStudy (anonymous):

indeed i did

OpenStudy (anonymous):

well i multiplied both sides by -2 to eliminate the denominator on the right

OpenStudy (anonymous):

ah ok, thanks